Cindy's age and her older cousin Robert's age together equal 28. Cindys age is 3 times less than Robert's. How old are the cousins?
Answer by stanbon(75887)

Cindy's age and her older cousin Robert's age together equal 28. Cindys age is 3 times less than Robert's. How old are the cousins?
Equations:
c + r = 28
c = r-3
----
Rearrange:
c + r = 28
c - r = -3
---
Add and solve for "c":
2c = 25
c = 12.5 (Cindy's age)
---
Solve for "r":
c + r = 28
12.5 + r = 28
r = 15.5 (Robert's age)
==================
Cheers,
Stan H.